OpenStudy (anonymous):

If sin A = 5/13, cos B = 3/5, and 0 < B < pi/2 < A < pi, what is cos (2A-B)?

OpenStudy (anonymous):

sin A = 5/13 cos A = -12/13 tan A = -5/12 sin B = 4/5 cos B = 3/5 tan B = 4/3

OpenStudy (anonymous):

cos 2A = (-12/13)^2 - (5/13)^2 = 119/169 sin 2A = (2)(5/13)(-12/13) = -120/169

OpenStudy (anonymous):

cos(a+b) = cos(a) cos(b) - sin(a) sin(b)

OpenStudy (anonymous):

cos (2A-B) = cos 2A cos B + sin 2A sin B

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

therefore cos (2A-B) = (119/169)(3/5) + (-120/169)(4/5) = -123/845

OpenStudy (anonymous):

can the answer be negative?

OpenStudy (anonymous):

definitely.... good job...:)

OpenStudy (anonymous):

thanks!